OpenStudy (anonymous):

What is the radius of a sphere if its surface area is 2,122.64 square inches? (Use 3.14 for π.) 11 in. 13 in. 121 in. 169 in.

OpenStudy (anonymous):

Well, I just used the previous formula that we had, and then I just went down the multiple choice list.

OpenStudy (anonymous):

4 times pi times a multiple choice option, so is it:\[4 \times \pi \times 11\]\[4 \times \pi \times 13\]\[4 \times \pi \times 121\]\[4 \times \pi \times169\]

OpenStudy (anonymous):

where did u get 11,13,121?

OpenStudy (anonymous):

They're multiple choice options...

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

\[13^{2}=169\]\[169\times3.14=530.66\]\[530.66\times4=?\]

OpenStudy (anonymous):

2122.64 @OrangeMaster

OpenStudy (anonymous):

Yes, so the answer is 13